OpenStudy (anonymous):

I will love u if u do dis :) plz guide me through it dont giv direct answers 。◕ ‿ ◕。 i need to know how to do it Which of the following rational numbers is equal to 3 point 1 with bar over 1? A 22 over 9 B 24 over 9 C 26 over 9 D 28 over 9

OpenStudy (anonymous):

:) lol

OpenStudy (anonymous):

no its repeating _ the "bar over the one" means its like 3.1

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

k :) thx

OpenStudy (anonymous):

\[3.1111....=3+.1+.01+.001+.....=3+\frac{ .1 }{ 1-.1 }=3+\frac{ .1 }{ 0.9 }=3+\frac{ 1 }{ 9 }=\frac{ 28 }{3 }\]

OpenStudy (anonymous):

¯\(°_o)/¯

OpenStudy (aric200):

28/3?

OpenStudy (anonymous):

yeah thats confusing nuntheless. ¯\(°_o)/¯

OpenStudy (anonymous):

i have solved above by finding the sum of an infinite geometric series.

OpenStudy (anonymous):

/(O-o)\ wt

OpenStudy (anonymous):

╥﹏╥ ima work on some other questions then, ill be back...

OpenStudy (anonymous):

here is an other method let x=3.111.... (1) 10x=31.111... (2) subtract (1) from (2) 10 x-x=31.111....-3.111... 9x=28 \[x=\frac{ 28 }{ 9 }\]

OpenStudy (anonymous):

okay. ;-;